# Balancing LoRA Performance and Efficiency with Simple Shard Sharing

### v10

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), effectively reduce the number of trainable parameters in Large Language Models (LLMs). However, as model scales continue to grow, the demand for computational resources remains a significant challenge. Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time). This paper introduces MiSS(Matrix Shard Sharing ), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism. MiSS leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment. This method constructs the low-rank update matrix through the replication of these shared, partitioned shards. We also propose a hardware-efficient and broadly applicable implementation for MiSS. Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate MiSS’s superiority. The results show that MiSS significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput. By effectively balancing expressive power and resource utilization, MiSS offers a compelling solution for efficiently adapting large-scale models.
@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) [16, 17, 25] is a prevalent methodology for adapting these models to specific downstream tasks. However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques [23] have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) [8] has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
The improvements of LoRA are mimicked in two streams: (1) Adaptability [6, 12, 2]: This refers to the celerity with which the method converges to an optimal or near-optimal state. The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$. Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants [7, 14, 20]. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed. Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. Certain solutions may improve the performance of finetuning, but makes the initialization process significantly more complex or increase the training cost, leading to a worse complexity. (2) Efficiency [11, 19, 22]: This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead. Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process. However, further simplification of the LoRA structure may result in an additional loss in performance. However, the critical balance between achieving highly expressive architectures and maintaining algorithmic efficiency has often received insufficient attention.
In this work, we systematically investigate the Pareto frontier of the trade-off between Adaptability and Efficiency in PEFT. We contend that an ideal PEFT method must adeptly reconcile these two crucial aspects. To this end, we first conduct a series of foundational experiments focusing on these dimensions, encompassing a simulated pre-training and fine-tuning pipeline, computational complexity analysis, and initialization time tests. Our analysis reveals that while LoRA-like methods leveraging shard sharing can achieve excellent computational efficiency, their expressive power may be constrained by such sharing, sometimes leading to performance below the standard LoRA baseline. Conversely, SVD-based modifications to the original LoRA often exhibit strong adaptability, yet their practical efficiency can be undermined by complex initialization procedures and more intricate forward and update mechanisms.
Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria. This approach, which we introduce as MiSS,  exhibits notable strengths. In terms of adaptability, MiSS demonstrates superior approximation fidelity compared to LoRA and several of its variants. Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability. Furthermore, concerning efficiency, MiSS benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts. The formulation of MiSS can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
@@
MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
We perform scaling experiments to verify our preliminary conclusions. Our extensive evaluation shows that MiSS demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method. Moreover, we present the training dynamics i.e., loss curves, which MiSS also showcase a significantly faster convergence rate compared to other baselines. In generation tasks, MiSS demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board. This demonstrates the feasibility of the framework. Furthermore, we validate the efficiency of MiSS  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
Our contributions can be summarized as threefolds:
@@
    In Section 3, we discuss the balancing Between Adaptability and Efficiency in LoRA-like methods with highly artificial controlled dataset to make the preliminary verification transparently and easy for replication.
@@
    In Section 4, we proposed MiSS, an efficient and adaptable structure with shard sharing mechanism.
  3. 3.
    In Section 5, we scale the experiments across a wide range of datasets and models, validated its effectiveness through extensive experiments.
def init(in_feats: int, out_feats: int):
weight = nn.Parameter(
torch.empty((out_features, in_features)) )
D = nn.Parameter(torch.zeros(self.r, out_features))
def forward(self, x):
result = F.linear(x, self.weight) # x: [B, T, C]
y = result + torch.sum(x.reshape
(*x.shape[:-1], x.size(-1)//self.r, self.r)\\
, dim=-2) @ self.miss
return y
Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
@@ 2 Preliminaries and Related Works
LoRA is the low rank adaptation towards full-param finetuning, and intuitively it downperforms than it. Several works propose diverse methods towards a better convergence and adaptability of LoRA. One compelling venue is to change the form of LoRA. PiSSA [14] optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), OLoRA [3] leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\bm{A}$ and $\bm{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning. LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency. DoRA [12] decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning. The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time.
@@ 3 No Free Launch: Balancing Between Adaptability and Efficiency
This section elucidates the fundamental trade-off inherent in LoRA-style PEFT techniques: the delicate balance between their adaptability and efficiency. Adaptability, in this context, refers to the capacity of a given method to emulate the performance benchmarks set by full-parameter fine-tuning. Conversely, efficiency encompasses the method’s judicious use of computational resources, specifically time and memory. We utilize highly artificial controlled dataset and model with relatively small parameter count to make the verification transparently and easy for replication.
We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning [13]. (2) LoRA [8]. (3) Alternatives to LoRA w/ different architectures, including: PiSSA [14], VeRA [11], DoRA [12] and MoRA [10]. (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA [19], MoS [22]. (1) An overview of their forward form, initialization method can be found at Table 1.
@@ Results.
Figure 2 illustrates the comparative adaptability of different methods. We utilize the minimum validation loss achieved by each approach as an indicator of its expressive capacity when approximating the performance of full-parameter fine-tuning. The results clearly demonstrate that methods leveraging singular value decomposition (SVD), such as PiSSA, attain a relatively low loss. Conversely, efficiency-focused techniques like MoS exhibit higher losses. A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power. Our method MiSS reaches a relatively advanced performance comparing to other variants.
@@ 4.1 Method Overview
In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{T}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$. This approach achieves efficiency by limiting the number of parameters. However, we observe that a repeating matrix—where a small matrix is replicated to form a larger one—can also be viewed as a low-rank structure. For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
@@
$(\operatorname{expand}(\bm{D}))^{\top}=\begin{bmatrix}(\mathbf{1}_{s_{1}}\bm{D_{1}})^{\top}\ (\mathbf{1}_{s_{2}}\bm{D_{2}})^{\top}\ \dots\ (\mathbf{1}_{s_{N}}\bm{D_{N}})^{\top}\end{bmatrix}$ (3)
@@ 4.2 Efficient Implementation of MiSS
$\Delta\bm{Wx}=\bm{D^{\top}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\top}S},\ \text{where}\ \bm{D}^{\top}\in\mathbb{R}^{d\times r}.$ (6)
Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\top}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{T}S$. Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension. E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index. This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.
(a) Loss-Token
(b) Loss-Time
Figure 3: Left. MiSS consumes less GPU memory and achieves higher throughput(# tokens per second). GPU memory usage refers to the amount of memory occupied by an RTX 4090 card. Right. The results of fine-tuning LLaMA2-7B on MetaMathQA.

### v11

# MiSS: Revisiting the Trade-off in LoRA with an Efficient Shard-Sharing Structure

@@ Abstract
Low-Rank Adaptation (LoRA) is a widely adopted technique for parameter-efficient fine-tuning, but its slow convergence has spurred the development of numerous variants. Nevertheless, existing methods often fail to improve performance, memory footprint, and computational efficiency simultaneously. To address this challenge, we revisit the causes of LoRA’s slow convergence. Building on these insights, we propose Matrix Shard Sharing (MiSS), which updates shards of the original weight matrix using a single shared trainable matrix $\bm{D}$, initialized to zeros. To simultaneously ensure computational efficiency, low memory footprint, and scalable serving, we introduce MiSSe. Both theoretical analysis and empirical results demonstrate that our method reduces optimization complexity without compromising performance, thereby achieving a more favorable trade-off among performance, memory, and efficiency. Furthermore, we conduct a comprehensive comparative analysis of various PEFT methods, evaluating their memory usage, initialization overhead, and computational efficiency. By mapping the Pareto frontier across these dimensions, we show that MiSS occupies a favorable position, effectively capturing the advantages of prior approaches.
https://github.com/Joluck/MiSS
@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) (Radford et al., 2019; Raffel et al., 2020; Yin et al., 2024) is a prevalent methodology for adapting these models to specific downstream tasks. However, full fine-tuning of all parameters is computationally prohibitive. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (Xu et al., 2023) have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (Ding et al., 2023; Liu et al., 2024; Biderman et al., 2024): This refers to the convergence speed at which the method reaches an optimal or near-optimal state. The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$. Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants (Hayou et al., 2024; Meng et al., 2024; Wang et al., 2024a). By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed. Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. However, many of these methods overlook issues of computational efficiency and overall training overhead. For example, PiSSA (Meng et al., 2024) requires a lengthy initialization process, while LoRA-GA (Wang et al., 2024b) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers. (2) Efficiency (Kopiczko et al., 2024; Wang et al., 2024c; 2025): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead. Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process. Although LoRA has demonstrated significant advantages in reducing parameter scale and computational cost, its effectiveness still falls short of fully matching full fine-tuning. To address this gap, researchers have proposed an increasing number of LoRA variants that gradually approach the performance of full fine-tuning. This raises a natural question:
Given the inherent challenge for LoRA and its variants to balance performance, memory, and efficiency, how can we achieve an effective trade-off among all three dimensions?
@@
MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\intercal}S}$ $\bm{D}\sim 0$
To strike a balance between performance, memory, and efficiency, we re-examined the key factors affecting LoRA’s slow convergence. Through an analysis of $S^{2}$FT (Yang et al., 2024) and LoRA+ (Hayou et al., 2024), we identified a critical phenomenon: During the LoRA fine-tuning process, both matrices $B$ and $A$ need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence. LoRA+ alleviates this issue by modifying the initial gradients, allowing the fine-tuning process to approximate full fine-tuning better. In contrast, $S^{2}$FT fixes one matrix as an orthogonal matrix, reducing the degrees of freedom in parameter updates and lowering optimization complexity, thereby enabling faster alignment with the optimal update direction. Inspired by these insights, we hypothesize that training only a single matrix could simplify optimization without sacrificing expressive capacity. We therefore propose Matrix Shard Sharing (MiSS), a method that updates a shard of the original weight matrix using a single, shared trainable matrix $\bm{D}$, initialized to zero. Thus, our approach maintains the low-rank property of the matrices while offering a more efficient alternative to $\bm{BA}$ updates in terms of computation.
Gradient Norm Analysis.
We analyze the initial gradient norm to verify our preliminary conclusions. In the experimental sections of the PiSSA, $S^{2}$FT, and LoRA-GA papers, we observed that LoRA exhibits a very small initial gradient norm compared to full fine-tuning, which shows a much larger one. Notably, all these improved methods share a common characteristic: their initial gradient norms are significantly larger than LoRA, and their early-stage convergence speed is comparable to that of full fine-tuning. Motivated by this, we evaluated the initial gradient norms of different methods across various models and datasets to examine whether MiSS follows the same pattern as other LoRA variants. The experimental results (Figure1,4) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence. This also supports the hypothesis that optimizing a single matrix is inherently simpler.
326412825602468Matrix rankGradientNorm
326412825601234Matrix rankGradientNorm
3264128256081624Matrix rankGradientNorm
32641282560481216Matrix rankGradientNormFinetuneLoRAPiSSAMiSS
Figure 1: Comparison of initial gradient norms across different training methods and the effect of rank. Results are shown for LLaMA2-7B and Qwen3-4B on the Math and Code datasets.
MiSSe
To achieve both computational efficiency and structural generality, we introduce an efficient and mathematically equivalent formulation of MiSS that offers improved time and space complexity. We further conduct extensive experiments (Table 5.3) to validate its effectiveness.
We first evaluate MiSS on both Natural Language Understanding (NLU) and Generation (NLG) tasks, assessing its performance and scalability. Our results show that MiSS consistently outperforms LoRA and its variants across diverse LLM architectures, establishing new state-of-the-art results on a wide range of metrics. We then analyze the Pareto frontier of the adaptability-efficiency trade-off in PEFT. We argue that an ideal PEFT method should effectively balance these two essential dimensions. To this end, we conduct a series of foundational experiments, including a simulated pre-training and fine-tuning pipeline, computational complexity analysis, and initialization time evaluation. With comprehensive empirical results, we demonstrate that MiSS achieves a favorable balance across three key dimensions performance, memory, and efficiency, highlighting its practicality as a general PEFT solution.
Our contributions can be summarized as follows:
@@
    We propose MiSS, an efficient and adaptable structure with a shard-sharing mechanism, striking an effective balance among three essential properties—performance, memory efficiency, and computational efficiency.
@@
    Through large-scale experiments across diverse datasets and model architectures, we provide a comprehensive evaluation of multiple PEFT methods. Our empirical results conclusively demonstrate that MiSS achieves a superior balance among these three properties compared to existing alternatives.
@@ 2 Preliminaries and Related Works
LoRA is the low rank adaptation towards full-param finetuning, and intuitively it downperforms than it. Several works propose diverse methods towards a better convergence and adaptability of LoRA. One compelling venue is to change the form of LoRA. PiSSA (Meng et al., 2024) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\bm{A}$ and $\bm{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning. LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency. DoRA (Liu et al., 2024) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning. The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time.
@@ 3 No Free Launch: Balancing Between Adaptability and Efficiency
This section elucidates the fundamental trade-off inherent in LoRA-style PEFT techniques: the delicate balance between their adaptability and efficiency. Adaptability, in this context, refers to the capacity of a given method to emulate the performance benchmarks set by full-parameter fine-tuning. Conversely, efficiency encompasses the method’s judicious use of computational resources, specifically time and memory. We utilize highly artificial controlled dataset and model with a relatively small parameter count to make the verification transparently and easy for replication.
We considered diverse methods  [^1] We have not included methods such as LoRA-GA (Wang et al., 2024b) or LoRA+ (Hayou et al., 2024) in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (Lv et al., 2024). (2) LoRA (Hu et al., 2021). (3) Alternatives to LoRA w/ different architectures, including: PiSSA (Meng et al., 2024), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024) and MoRA (Jiang et al., 2024). (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (Wang et al., 2024c), MoS (Wang et al., 2025). (1) An overview of their forward form, initialization method can be found at Table 1.
@@ Results.
Figure 2 illustrates the comparative adaptability of different methods. We utilize the minimum validation loss achieved by each approach as an indicator of its expressive capacity when approximating the performance of full-parameter fine-tuning. The results clearly demonstrate that methods leveraging singular value decomposition (SVD), such as PiSSA, attain a relatively low loss. Conversely, efficiency-focused techniques like MoS exhibit higher losses. A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components, which may inherently constrain their expressive power. Our method MiSS reaches a relatively advanced performance comparing to other variants.
@@ 4.1 Method Overview
In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{\intercal}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$. This approach achieves efficiency by limiting the number of parameters. However, we observe that a repeating matrix—where a small matrix is replicated to form a larger one—can also be viewed as a low-rank structure. For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
def init(in_features: int, in_features: int, rank: int):
self.r = rank
self.weight = $n$.Parameter(torch.empty((out_features, in_features)))
self.D = $n$.Parameter(torch.zeros(self.r, out_features))
def forward(self, x):
result = F.linear(x, self.weight) # x: [B, T, C]
y = result + x @ self.D.expand(in_features//self.r,1)
return y
Figure 3: Left. Structural diagram of $\Delta\bm{W}$ in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
@@
$(\operatorname{expand}(\bm{D}))^{\intercal}=\begin{bmatrix}(\mathbf{1}_{s_{1}}\bm{D_{1}})^{\intercal}\ (\mathbf{1}_{s_{2}}\bm{D_{2}})^{\intercal}\ \dots\ (\mathbf{1}_{s_{N}}\bm{D_{N}})^{\intercal}\end{bmatrix}$ (3)
@@ 4.2 Efficient Implementation of MiSS
$\Delta\bm{Wx}=\bm{D^{\intercal}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\intercal}S},\ \text{where}\ \bm{D}^{\intercal}\in\mathbb{R}^{d\times r}.$ (6)
Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\intercal}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{\intercal}S$. Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension. E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index. This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.3.